Circular No. 8448

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 WR_{9} (LINEAR) An apparently asteroidal object discovered by LINEAR, and published

as 2004 WR_{9} on MPO 120656 (discovery observation below), has been found to show cometary appearance. P. Birtwhistle, Great Shefford, U.K., reports that CCD images taken on Dec. 7.0 UT with a 0.30-m f/6.3 reflector show the object as having a tail 20′′ long in p.a. 235^{o} and a very concentrated coma with a diameter of 9′′, while on images from Dec. 3.0, there had been a suggestion of a 10′′ tail in p.a. 250^{o}. C. Hergenrother reports that R-band images taken with the Catalina 1.54-m reflector on Dec. 8.45 show a highly condensed coma 11′′ in diameter and a faint tail > 25′′ long in p.a. 235^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Nov. 22.32392  5^{h}31^{m}41.^{s}94  +30^{o}19′19′′.2  18.7

The available astrometry, the following orbital elements, and an ephemeris appear on MPEC 2004-X28.

T = 2005 Jan. 11.3666 TT  ω = 70.^{o}5784
e = 0.683440  Ω = 25.4444 }2000.0
q = 1.917200 AU  i = 5.0473
a = 6.056348 AU n^{o} = 0.0661284 P = 14.90 years

SUPERNOVAE 2004gn AND 2004go Further to IAUC 8447, W. Li reports the LOSS discovery of an appar-

ent supernova (mag ∼ 16.6) on an unfiltered KAIT image taken on Dec. 1.57 UT. SN 2004gn is located at α = 12^{h}34^{m}12^{s}.10, Decl. = +2^{o}39′34′′.4 (equinox 2000.0), which is 55′′.0 east and 21′′.4 north of the nucleus of NGC 4527. A KAIT image on June 10.21 showed nothing at this position (limiting mag ∼ 19.0). H. Yamaoka, Kyushu University, reports that K. Itagaki, Yamagata, Japan, has confirmed the existence of the new object at red mag ∼ 16.7 in unfiltered CCD images taken on Dec. 2.806 (the position end figures given as 12^{s}.10, 34′′.7).

H. Pugh and Li report the LOSS discovery of another apparent supernova on unfiltered KAIT images taken on Nov. 18.32 (at mag 16.5), 23.25 (mag 16.6), Dec. 4.32 (mag 17.3), and 5.29 UT (mag 17.3). SN 2004go is located at α = 2^{h}55^{m}41^{s}.76, δ = −14^{o}14′14′′.5 (equinox 2000.0), which is 34′′.8 west and 105′′.7 south of the nucleus of IC 270. A KAIT image taken on Oct. 14.43 showed nothing at this position (limiting mag ∼ 19.0).

2004 December 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green